# Physics-Informed Detection of Friction Anomalies in Satellite Reaction Wheels

### v1

@@ Abstract
As the number of satellites in orbit has increased exponentially in recent years, ensuring their correct functionality has started to require automated methods to decrease human workload. In this work, we present an algorithm that analyzes the on-board data related to friction from the Reaction Wheel Assemblies (RWA) of a satellite and determines their operating status, distinguishing between nominal status and several possible anomalies that require preventive measures to be taken. The algorithm first uses a model based on hybrid systems theory to extract the information relevant to the problem. The extraction process combines techniques in changepoint detection, dynamic programming, and maximum likelihood in a structured way. A classifier then uses the extracted information to determine the status of the RWA. This last classifier has been previously trained with a labelled dataset produced by a high-fidelity simulator, comprised for the most part of nominal data. The final algorithm combines model-based and data-based approaches to obtain satisfactory results with an accuracy around $\mathbf{95\boldsymbol{\%}}$.
Index Terms:
satellite, reaction wheel assembly, friction, anomaly detection, physics-informed
@@ I Introduction
As satellites orbiting Earth have become more frequent during the last decades [1], the need to maintain their correct functioning has become a major concern of space agencies around the world. Being in space during their whole lifetime, satellite repairs are, in general, unfeasible, so failures usually lead to the termination of the mission. Even when a repair is possible, it tends to be very expensive, as was the case of the installation of a new camera for the Hubble telescope lens, which required an entire mission [2]. This makes the capability of detecting impending failures a fundamental component of their monitoring.
@@
Among the components that can fail in a satellite, the Reaction Wheel Assemblies (RWAs) are of particular interest. RWAs are a fundamental component of the Attitude and Orbit Control System (AOCS) of the satellite. Their purpose is to exert a torque on the body of the satellite to control its attitude. Rather than using action-reaction principles as a Reaction Control System (RCS) does, an RWA uses angular momentum conservation to achieve greater accuracy with no propellant consumption, which is highly advantageous for star tracking and long-duration missions.
@@
The operating principle of RWAs is fairly simple. Inside the assembly, a flywheel is kept rotating permanently. When there is a need to rotate the satellite in one direction, the disc is accelerated in the opposite direction using an electric motor. Due to the conservation of angular momentum in space, this produces the desired torque in the main body of the satellite [3].
Figure 1: Cross-section of a Reaction Wheel Assembly, adapted from [4].
@@
To get complete control of the three axes of rotation of the satellite, at least 3 RWAs need to be mounted in a satellite. However, to achieve redundancy and some flexibility in the speed of rotation of the discs, it is common to install 4 RWAs instead [5].
@@
The RWA is, then, a mechanical component in almost permanent motion, making it more prone to failure than other fixed satellite parts. Mechanical degradation, vibrations, and wear can occur in this element, compromising its integrity [6]. Particularly, the bearings, where contact between the rotating disc and the non-moving satellite body takes place, are home to several complex phenomena involving elastohydrodynamics (interaction between a moving fluid, such as lubricant, and deformable solid material). These phenomena cause sudden changes in friction at the bearings that make the wheel modeling particularly challenging. A failure of a RWA might imply losing all device functionality and can seriously compromise the mission’s success. Such has been the case in space missions Kepler [7], Dawn [8], and SDO [9], in which the capabilities of the satellites were compromised after RWA failures due to the loss of authority over the torque of the satellite [10, 9].
@@
Although there have been attempts at explaining the causes of such failures [11], the complexity in the dynamics of the bearings makes it hard to determine the condition of the component, and the limited number of sensors in the RWA implies that a complete monitoring of the situation is impossible. Even though the nature of these failures remains unknown, recent studies have shown that they are usually preceded by changes in the behavior of the friction at the bearings, which can be observed using on-board sensors commonly installed in RWAs [12]. Such changes of behavior (that is, anomalies) are usually caused by changes in the properties of the RWA, such as alterations of the lubricant or an increase in the contact surface in the bearings.
It is then of interest to detect these changes in the behavior, as they can be used to anticipate possible future failures of the wheel and/or take mitigative measures to improve the health of the RWA [4]. However, not all changes imply a future failure, so it is also necessary to distinguish between different types of anomalous behavior to provide the appropriate warnings. Therefore, for this work, it is not enough to perform anomaly detection, but categorization should also be considered.
@@
In this paper, we develop a method that analyzes the on-board data collected from an RWA during a window of time of a few hours and determines whether an anomaly is present in the RWA and, if positive, which one among those known, with the entire process carried out in the satellite thanks to its low computation cost. We propose the following contributions:
@@
  • We derive a novel model for friction systems that incorporates sudden variations of friction due to quick changes in the internal geometry of the system. The model can be implemented without knowledge of the actual causes of these phenomena, but rather just their statistical properties, which can be obtained experimentally.
@@
  • We provide an end-to-end algorithm that takes a window of a few hours of measurements by on-board sensors to determine whether the observed friction characteristics of the RWA belong to nominal or anomalous conditions. The method combines the use of the model together with data-driven approaches to provide an accurate diagnosis.
  • We include in the approach a modification to the classification of changepoints that accounts for the possibility of false negatives. This makes the algorithm robust to incorrectly detected changepoints, an unavoidable phenomenon due to the stochastic nature of the signal measured.
@@
  • We thoroughly validate the complete algorithm and each of its components using high-fidelity simulation data from Thales Alenia Space (TAS), with no knowledge of the model used to generate it.
Literature review
@@
Anomaly and/or fault detection in satellites has received considerable interest in recent years, primarily due to their proliferation and the need for constant monitoring. Research done in this direction has attempted to produce algorithms capable of analyzing all the information coming from the satellite to determine a wide variety of possibly faulty conditions. The high availability of data and the lack of models for the complete satellite has led to almost exclusively data-driven approaches, using tools such as Support Vector Machines (SVM) [13], Relevance Vector Machines (RVM) [14], Principal Component Analysis (PCA) [15], or LSTM networks [16]. These ideas have led to algorithms used in actual space missions, like Orca [17] in the International Space Station and the IMS tool [18] in the Space shuttle.
@@
The problem of fault and anomaly detection for RWAs has been analyzed mainly through machine learning [21, 22]. Although these methods have been effective in their purpose, they consider the behavior of the electric motor rather than focusing on the friction torque, so they are not so precise in determining abnormal behavior of the latter. This was done only recently using unsupervised machine learning [23], but considering only normal and abnormal behavior, that is, without isolation of the anomaly.
Closely related to RWAs, rotary machinery analysis is a richer field of literature. Although purely model-based approaches have been proposed in the past [24], they are usually focused on alterations in the vibration patterns of the machine. Most of the research is based on data-driven techniques, capable of assessing a wider variety of faulty conditions usually related to problems in the bearings, as highlighted in the survey [25].
Most of the rotary machinery analysis techniques follow a common approach: first, some signals are measured from the machine (usually displacement or acceleration), then features are extracted (statistical ones like mean and variance, or frequency-related like Power Spectral Density), and finally, some machine learning methods are used to determine the status of the machine from these features. Such methods include nearest neighbor [26], neural networks [27], SVM [28], linear regressions [29], or Broad Learning Systems (BLS) [30], among others.
These methods have achieved important improvements in recent years by incorporating knowledge of a model of the machine, leading to the family of methods known as hybrid methods [31]. A popular technique is to fuse these two sources of knowledge through Physics Informed Machine Learning (PIML), which has experienced a rapid development in the last years [32].
Although evident progress has been achieved as new features and classification methods have been developed, their focus on the vibrational behavior of the machine makes them unsuitable for satellite applications, where this behavior is not of much interest. Regarding friction-related research, most of the interest is on the low-speed behavior, where complex phenomena like the Stribeck curve or stick-slip occur [33, 34, 35]. Research on switching behavior in friction has also experienced some attention, but it is restricted to deterministic systems[36].
@@ II Problem Formulation
We study the problem of determining if the friction torque of an RWA has an anomalous behavior based on a window of measurements taken by its sensor system at fixed time intervals. Each measurement at time $t_{k}$ (where $t\in\mathbb{R}$, and $k\in\mathbb{N}$ is the index of the measurement) consists of the spin rate $\omega_{k}$ of the RWA, the voltage $V_{k}$, and the current $I_{k}$ in the motor that accelerates (or deccelerates) the RWA. Using conservation of angular momentum, energy equations at the motor and filtering techniques, we can obtain an unbiased estimation of the friction torque $f_{k}$ at the bearings of the RWA using an estimator:
$J\dot{\omega}$ $=T_{\text{motor}}+f=K_{T}I+f$ (1a)
$\hat{f}_{k}$ $=J\frac{\omega_{k+1}-\omega_{k-1}}{t_{k+1}-t_{k-1}}-K_{T}I_{k}$ (1b)
where $T_{\text{motor}}$ is the torque produced by the motor, $J$ is the wheel inertia and $K_{T}$ is the torque constant of the motor. Additionally, the estimated $\hat{f}_{k}$ can be filtered to remove the effect of noise. In this paper, we do not consider the details of this estimation process, but rather assume that $\hat{f}_{k}=f_{k}+v_{k}$, where $v_{k}$ is random noise with zero mean, assuming the estimation is unbiased.
Since we are interested only in the friction characteristics of the RWA, the input to our algorithm is a sequence of pairs of spin rate and estimated friction torque obtained in a window of time:
@@
Both the spin rate and the estimated friction torque are normalized using average values in historical data, so the variables take values of order $1$. To ease the reading of the paper, a table with all the indices used in the paper is given in Table I. Subscripts and superscripts that are part of the name of the variable are written upright (like $N_{\mathrm{cp}}$) and those that are indices are written in italics (like $\omega_{k}$).
Symbol Meaning Range
@@
TABLE I: Indexes used in the paper
In this paper, we want to detect anomalies, defined as abnormal behavior in the friction torque of the RWA. Several possible anomalies are considered, each due to different phenomena occurring in the RWA. For a window of measurements $X$, any anomaly can be active or inactive: the status of the RWA is defined as a vector of boolean values specifying whether each anomaly is active or inactive: $\Theta=\{0,1\}^{N_{\mathrm{anom}}}$, where $N_{\mathrm{c}}$ is the number of anomalies considered.
Anomalies do not have an explicit mathematical description. Instead, our knowledge about them comes from a labeled dataset $D$ with windows of measurements annotated with the anomalies affecting the RWA that produced such measurements:
$D=\{(X_{\ell},\theta_{\ell})\}_{\ell=1}^{N_{\mathrm{D}}},$ (3)
where $\theta_{\ell}\in\Theta$. This dataset can be obtained from ground tests, in-orbit experience, or hight-fidelity simulators, for example.
To illustrate the shape of the datapoints in such a dataset, and how they might differ depending on the anomaly, we show three examples taken from the TAS dataset in Figure 2. Some early observations indicate that, even in the nominal case, there are sudden variations of the friction independent of the spin rate (for example, the one close to measurement $4000$ in the nominal case). Sudden changes in the bearings cause such variations due to, for example, oil accumulations in small pockets. Although their presence does not necessarily imply an anomaly in the RWA, they must be considered, since excessive friction changes due to these variations are usually concerning. A model that captures the behavior of an RWA must, therefore, model these sudden friction changes.
Regarding anomalies, Anomaly 1 produces a constant increase of friction torque at all spin rates, while Anomaly 2 increases friction torque proportionally with the spin rate. Additional anomalies relate to friction differences produced in the sudden events described in the previous paragraph.
@@
Based on these observations and the given dataset from TAS, we state the following problem.
Problem Statement: Create a model that captures the behavior observed in the TAS dataset, and use it to produce an algorithm that can determine the anomaly affecting an RWA based on a window of spin rate and friction torque measurements.
@@ III RWA Modeling
To solve the problem formulated in the previous section, we propose a hybrid approach to overcome the usual limitation of purely data-based and model-based methods. In this section, we derive the model used for the model-based part, developed from a combination of dataset observation, knowledge of the RWA mechanism, and literature on the topic.
@@ III-A Friction Model
Since the data available for the detection of anomalies consists of spin rate and friction torque signals, we discard the dynamical part shown in (1) and model the RWA as a friction system, that is, an algebraic relation between friction and spin rate:
$f_{k}$ $=f_{k}^{\mathrm{d}}\text{sign}(\omega_{k})+f^{\mathrm{v}}\omega_{k}+v_{k}$ (4a)
$f_{k}^{\mathrm{d}}$ $=f^{\bar{\mathrm{d}}}+\sum_{s=1}^{N_{\textrm{s}}}f_{k}^{s}$ (4b)
@@
where (4a) corresponds to the classical dry and viscous friction with additive Gaussian noise $v$. The dry and viscous friction coefficients $f_{k}^{\mathrm{d}}$ and $f^{\mathrm{v}}$, respectively, are at the basis of any friction model [37], but note here that the dry coefficient is not constant and can change with time.
@@
Each component of (4a) models a different friction factor. Starting from the end, the Gaussian noise $v$ has three sources: randomness in the motion of the bearing and the contact inside it, sensor noise, and inaccuracies in the friction torque estimation. Among all possible noise models available, the Gaussian one was chosen for its theoretical properties and its popularity in the literature [38, 39, 40].
@@
The viscous friction, on the other hand, is produced by the lubricant inside the bearing [41]. This parameter is known to be affected by the temperature and the pressure of the lubricant, as well as its age [42]. However, these variations take place in longer time intervals than we analyze; therefore, we can assume this coefficient to be constant during the periods of time analyzed with the algorithm. Note that, although constant, this coefficient can still be affected by an anomaly that increases its value.
Finally, the dry friction models the contact between solid surfaces. In the bearings of an RWA, this contact occurs between the the inner and outer races of the bearing and the balls, and between the balls themselves [20]. The contact surface at every instant is not necessarily constant, which can lead to changes in the dry friction coefficient. To represent these changes, we first define in (4b) the dry friction coefficient as the sum of several terms, the first of which is the constant baseline dry friction $f^{\bar{d}}$ that remains constant at all times.
@@
The remaining components $f_{k}^{s}$ are variable in time, and capture the effect of variations in the contact surface of the bearings. The evolution in time of each of these components is described by what we refer to as a Friction Switching System (FSS). The motivation behind using several FSS in the model of a single RWA is to capture independent phenomena with different properties. Such phenomena can range from concentrations of pockets of oil (well reported in the literature [41] to displacements of the different components of the bearing (observed in the past but not well characterized so far [6]), which may have very different time scales. By modeling them with different FSS, we can also distinguish anomalies affecting one or another phenomenon, enhancing the accuracy of the categorization of anomalies. An example of using several FSS in a single RWA model is given in Example 1.
III-B Friction Switching System
We define a Friction Switching System (FSS) as a hidden Markov Model (HMM) [43] with state $(q,f,\tau)$ where $q\in\llbracket 1,q_{\text{max}}\rrbracket$, $f\in\mathbb{R}$ and $\tau\in\mathbb{N}^{\geq 0}$, where $\mathbb{R}$ is the set of real numbers and $\mathbb{N}^{\geq 0}$ is the set of natural numbers including 0.

### v2

@@ Abstract
As the number of satellites in orbit has increased exponentially in recent years, ensuring their correct functionality has started to require automated methods to decrease human workload. In this paper, we present an algorithm that analyzes the on-board data related to friction from the Reaction Wheel Assemblies (RWAs) of a satellite and determines their operating status, distinguishing between nominal status and several possible anomalies that require preventive measures to be taken. The algorithm is based on a hybrid system model and combines changepoint detection, dynamic programming, and maximum likelihood estimation. A classifier then uses the extracted information to determine the status of the RWA. The classifier is trained on a labelled dataset produced by a high-fidelity simulator, comprised for the most part of nominal data. The algorithm combines model-based and data-based approaches to obtain satisfactory results with an accuracy around $\mathbf{95\boldsymbol{\%}}$, a false positive rate below $\mathbf{3\boldsymbol{\%}}$, and an execution time in the order of a tenth of a second, making it suitable to run on-board. Moreover, the algorithm allows for easy tuning and explainability of the results obtained.
@@ I-A Motivation
As satellites orbiting Earth have become more frequent during the last decades [34], the need to maintain their correct functioning has become a major concern of space agencies. Often, satellite repairs are unfeasible, so failures usually lead to the termination of the mission. Even when repair is possible, it tends to be very expensive, as was the case of the installation of a new camera for the Hubble telescope lens, which required an entire mission [11]. This makes the capability of detecting impending new failures a fundamental component of any satellite system.
@@
Among the components that can fail in a satellite, the Reaction Wheel Assemblies (RWAs) are of particular interest. RWAs are fundamental components of the Attitude and Orbit Control System (AOCS). Their purpose is to exert a torque on the body of the satellite to control its attitude. Rather than using action–reaction principles as a Reaction Control System (RCS) does, an RWA uses angular momentum conservation to achieve greater accuracy with no propellant consumption, which is highly advantageous for star tracking and long-duration missions.
@@
The operating principle of RWAs is fairly simple. Inside the assembly, a flywheel is kept rotating permanently, see Figure 1. When there is a need to rotate the satellite in one direction, the flywheel is accelerated in the opposite direction using an electric motor. Due to the conservation of angular momentum, this produces the desired torque in the main body of the satellite [37].
Figure 1: Cross-section of an RWA, adapted from [38].
@@
To control the three axes of rotation of the satellite, at least three RWAs need to be mounted. However, to achieve redundancy and some flexibility in the speed of rotation of the flywheels, it is common to install four RWAs[7].
@@
As an RWA is a mechanical component in almost permanent motion, it is more prone to failure than other satellite parts. Mechanical degradation, vibrations, and wear occur, compromising its integrity [24]. Particularly, the bearings, where contact between the rotating disc and the non-moving satellite body takes place, are home to several complex phenomena involving elastohydrodynamics (interaction between a moving fluid, such as lubricant, and a deformable solid material). These phenomena cause sudden changes in friction at the bearings that make the wheel modeling particularly challenging. A failure of an RWA might imply losing satellite functionality and can seriously compromise the mission. This was the case in space missions Kepler [25], Dawn [40], and SDO [9], in which the capabilities of the satellites were compromised after RWA failures [43, 9].
@@
Although there have been attempts at explaining the causes of RWA failures [6], the complexity of the bearing inner dynamics makes it hard to determine the precise condition of the component. Additionally, limited number of RWA sensors implies that a complete monitoring of the situation is impossible. Even though the nature of these failures remains unknown, recent studies have shown that they are usually preceded by changes in the behavior of the friction at the bearings, which can be observed using on-board sensors [23]. Such anomalies are usually caused by RWA changes, such as alterations of the lubricant or an increase in the contact surface in the bearings. It is of interest to detect these changes as they can be used to anticipate possible future failures of the wheel and take measures to improve the health of the RWA [38]. However, not all changes imply a future failure, so it is necessary to distinguish between different types of anomalous behavior to provide the appropriate warnings. Therefore, it is not enough to perform anomaly detection, but categorization should also be considered.
I-B Contribution
@@
In this paper, we develop a method that analyzes the on-board data collected from an RWA during a window of time of a few hours to determine whether anomalies are present and, if positive, which ones. The algorithm can be implemented onboard the satellite thanks to its low computation cost. We make the following specific contributions:
@@
  • We derive a novel friction model that incorporates sudden friction variations due to quick changes in the internal geometry of the system. The model represents their statistical properties, which can be obtained experimentally.
@@
  • An end-to-end friction detection algorithm is derived that takes a window of a few hours of measurements by on-board sensors to determine whether the observed friction characteristics of the RWA belong to nominal or anomalous conditions. The method combines physical model knowledge with data to provide an accurate diagnosis procedure.
  • A modification in the changepoint classification algorithm that accounts for the possibility of false negatives, making the algorithm robust to false positives in the detection of changepoints.
@@
  • The algorithm and each of its components are experimentally evaluated using high-fidelity simulation data from Thales Alenia Space (TAS)  [^1] https://www.thalesaleniaspace.com, with no prior knowledge of the simulation model.
I-C Related work
@@
Anomaly and fault detection in satellites has received considerable interest in recent years, primarily due to their proliferation and the need for constant monitoring. Research done in this direction has attempted to produce algorithms capable of analyzing all the information coming from the satellite to determine a wide variety of possibly faulty conditions. The high availability of data and the lack of models for the complete satellite has led to almost exclusively data-driven approaches, using tools such as Support Vector Machines (SVM) [12], Relevance Vector Machines (RVM) [13], Principal Component Analysis (PCA) [14], or LSTM networks [21]. These ideas have led to algorithms used in actual space missions, like Orca [3] in the International Space Station and the IMS tool [22] in the Space shuttle.
@@
The problem of fault and anomaly detection for RWAs has been analyzed mainly through machine learning [4, 1]. Although these methods have been effective in their purpose, they consider the behavior of the electric motor rather than focusing on the friction torque, so they are not so precise in determining abnormal behavior of the latter. This was done only recently using unsupervised machine learning [36], but considering only normal and abnormal behavior, that is, without isolation of the anomaly. Although research have been done considering different possible anomalies, their are limited to increases in either dry or viscous friction [41].
Closely related to RWAs, rotary machinery analysis is a richer field of literature. Although purely model-based approaches have been proposed in the past [31], they are usually focused on alterations in the vibration patterns of the machine. Most of the research is based on data-driven techniques, capable of assessing a wider variety of faulty conditions usually related to problems in the bearings [19].
Most of the rotary machinery analysis techniques follow a common approach: first, some signals are measured from the machine (usually displacement or acceleration), then features are extracted (statistical ones like mean and variance, or frequency-related like Power Spectral Density), and finally, some machine learning methods are used to determine the status of the machine. Such methods include nearest neighbor [35], neural networks [28], SVM [46], linear regressions [42], or Broad Learning Systems (BLS) [29], among others.
These methods have achieved important improvements in recent years by incorporating knowledge of a model of the machine, leading what is known as hybrid methods [27]. A popular technique is to fuse these two sources of knowledge through Physics Informed Machine Learning (PIML), which has experienced a rapid development in the last years [45].
Although evident progress has been achieved as new features and classification methods have been developed, their focus on the vibrational behavior of the machine makes them unsuitable for satellites, where this behavior is not of much interest. Regarding friction-related research, most of the interest is on the low-speed behavior, where complex phenomena like the Stribeck curve or stick-slip occur [39, 15]. Research on switching behavior in friction has also received some attention, but restricted to deterministic systems[32].
@@
I-D Outline
The paper is structured as follows: Section II states the problem. In Section III, we introduce the RWA model used for the algorithm design. Section IV develops the algorithm. In Section V, we comment on the tuning of the algorithm’s parameters. Section VI describes the experimental evaluation and illustrates the performance of the algorithm.. Finally, we provide conclusions and possible future work in Section VII.
@@ II Problem Formulation
We study the problem of determining if the friction torque of an RWA has an anomalous behavior based on a sequence of measurements taken by its sensor system at fixed time intervals. Each measurement at time $t_{k}\in\mathbb{R}$, where $k\in\mathbb{N}$ is the index of the measurement, consists of the spin rate $\omega_{k}$ of the RWA, the voltage $V_{k}$, and the current $I_{k}$ in the motor that accelerates (or deccelerates) the RWA. Using conservation of angular momentum, energy equations at the motor, and filtering techniques, we obtain an unbiased estimation of the friction torque $f_{k}$ at the bearings of the RWA using the estimator
$\hat{f}_{k}$ $=J\frac{\omega_{k+1}-\omega_{k-1}}{t_{k+1}-t_{k-1}}-K_{T}I_{k},$ (1)
which follows from
$J\dot{\omega}$ $=T_{\text{motor}}+f=K_{T}I+f$ (2)
where $T_{\text{motor}}$ is the torque produced by the motor, $J$ is the wheel inertia, and $K_{T}$ is the motor torque constant. The estimated $\hat{f}_{k}$ can be filtered to remove the effect of noise. In this paper, we do not consider the details of this filtering process, but rather assume that $\hat{f}_{k}=f_{k}+v_{k}$, where $v_{k}$ is random noise with zero mean, as the filter is expected to be unbiased.
Since we are interested only in the friction characteristics of the RWA, the input to our algorithm is a sequence of spin rate and estimated friction torque pairs:
@@
This sequence is normalized using average values of historical data, so the variables take values of order one. All indices used in the paper are given in Table I. Subscripts and superscripts that are part of the name of the variable are written upright (like $N_{\mathrm{cp}}$) and those that are indices are written in italics (like $\omega_{k}$).
Index Meaning Range
@@
TABLE I: Indexes and their range
Consider we have $N_{D}$ sequences of measurements $X_{\ell}$, indexed by $\ell=[1,N_{D}]$. Each of these sequences is labeled with an RWA status $\theta_{\ell}$. We can collect these sequences and their corresponding labels into a dataset $D=\{(X_{\ell},\theta_{\ell})\}_{\ell=1}^{N_{\mathrm{D}}}$. This dataset can be obtained from ground tests, in-orbit measurements or, as in our cause, high-fidelity simulators.
The status of a RWA is related to its friction properties. In the dataset $D$, each sequence $X_{\ell}$ is labeled by the status $\theta_{\ell}$ of the RWA from which that sequence of measurements was taken. An RWA can be affected by $N_{\mathrm{an}}$ different anomalies, with two rules: anomalies are either active or inactive, and any number of anomalies can be active in an RWA simulteneously. This is modeled by defining the set of possible status as $\Theta=\{0,1\}^{N_{\mathrm{an}}}$, such that each element of $\theta$ indicates whether the corresponding anomaly is inactive (0) or active ($1$). The objective of the algorithm is, then, to determine for a new sequence $X$ what is the status of the RWA that generated it, that is, what anomalies affect it.
We evaluate our approach on data generated by a high-fidelity simulator developed by TAS, similar to the one used in [10]. To illustrate how this dataset looks like depending on the anomaly, we show three examples in Figure 2. Some early observations indicate that, even in the nominal case, there are sudden variations of the friction independent of the spin rate (for example, the one close to measurement $4000$ in the nominal case). Sudden changes in the bearings cause such variations due to, for example, oil accumulations in small pockets. Although their presence does not necessarily imply RWA anomalies, they must be considered, since excessive friction changes due to these variations are usually concerning. A model that captures the behavior of an RWA must, therefore, include these sudden friction changes.
Regarding differences between anomalies, Anomaly 1 produces a constant increase of friction torque at all spin rates, while Anomaly 2 increases friction torque proportionally with the spin rate. Additional anomalies relate to friction differences produced in the sudden events described in the previous paragraph.
@@
Problem Statement: Develop an algorithm that can detect the status $\theta$ of an RWA based on a sequence of measurements $X$ of spin rate and estimated friction torque, and train it based on historical labeled data given as a dataset $D$.
@@ III RWA Modeling
To solve the problem formulated in the previous section, we propose a hybrid approach combining data-based and model-based detection methods. In this section, we derive the model used for the model-based part, developed from a combination of dataset observation, knowledge of the RWA mechanism, and literature on the topic.
@@ III-A Friction Model
We model the RWA as a friction system:
$f_{k}=f_{k}^{\mathrm{d}}\text{sign}(\omega_{k})+f^{\mathrm{v}}\omega_{k}+v_{k}$ (4)
@@
where (4) corresponds to a classical dry and viscous friction model with additive Gaussian noise $v$ [2]. The dry and viscous friction coefficients are denoted $f_{k}^{\mathrm{d}}$ and $f^{\mathrm{v}}$, respectively, where the dry coefficient can change with time.
@@
Each component of (4) models a different friction factor. Starting from the end, the Gaussian noise $v$ has three sources: randomness in the motion of the bearing and the contact inside it, sensor noise, and inaccuracies in the friction torque estimation. Among all possible noise models available, the Gaussian one was chosen for its theoretical properties and its popularity in the literature [16, 8, 18], as well as for its good fit with the observed data.
@@
The viscous friction, on the other hand, is produced by the lubricant inside the bearing [26]. This parameter is known to be affected by the temperature and the pressure of the lubricant, as well as its age [44]. However, these variations take place over longer time intervals than we analyze; so we can assume this coefficient to be constant for our algorithm. Note that, although constant, this coefficient can still be affected by an anomaly that increases its value permanently.
Finally, the dry friction models the contact between solid surfaces, and is given by
$f_{k}^{\mathrm{d}}=f^{\bar{\mathrm{d}}}+\sum_{s=1}^{N_{\textrm{s}}}f_{k}^{s}.$ (5)
@@
In the RWA bearings, this contact occurs between the the inner and outer races of the bearing and the balls, and between the balls themselves [30]. The contact surface at every instant is not necessarily constant, which can lead to changes in the dry friction coefficient. To represent these changes, we define in (5) the dry friction coefficient as the sum of several terms, the first of which is the constant baseline dry friction $f^{\bar{d}}$ that remains constant at all times: The remaining components $f_{k}^{s}$ capture the effect of variations in the contact surface of the bearings. The evolution of each of these components is described by what we refer to as a Friction Switching System (FSS). The motivation behind using several FSS in the model of a single RWA is to capture independent phenomena with different properties. Such phenomena can range from concentrations of pockets of oil (well reported in the literature [26] to displacements of the different components of the bearing (observed in the past but not well characterized so far [24]), which may have very different time scales. By modeling them with different FSS, we can also distinguish anomalies affecting one or another phenomenon, enhancing the accuracy of the categorization of anomalies. An example of using several FSS in a single RWA model is given in Example 1.
III-B FSS
We define an FSS as a Hidden semi-Markov Model (HSMM) [48] with state $q\in\left[1,q_{\text{max}}\right]\subset\mathbb{N}^{\geq 0}$, the latter being the set of natural numbers including 0. An HSMM is an extension of the standard HMM, in which a duration $\tau\in\mathbb{N}^{\geq 0}$ is associated to the state of the system. The duration indicates the number of steps the HSMM stays in that state before the next transition. Every time there is a state transition, a new duration $\tau^{\prime}$ is drawn from a random distribution.
To formally define the HSMM we need a transition function and an observation function. The transition functions specifies how likely the system is to transition to state $q$ for a duration $\tau^{\prime}$ after having spent $\tau$ steps in state $q$. It is defined by the conditional probability function $P(q^{\prime},\tau^{\prime}|q,\tau)$.
For the observation function, we propose a simpler alternative to the standard HSMM formulation: when a transition takes place at step $k_{\mathrm{tr}}$, a friction value $f_{k_{\mathrm{tr}}}$ is drawn from a probability function given by $P^{\mathrm{f}}(f_{k_{\mathrm{tr}}}|q)$. Between transitions, the observed friction value produced by the FSS is constant, and equal to the friction $f$ generated in the last transition, that is, $f_{k\in[k_{\mathrm{tr}},k_{\mathrm{tr}}+\tau)}=f_{k_{\mathrm{tr}}}$.